\begin{thebibliography}{43}
\providecommand{\natexlab}[1]{#1}

\bibitem[{Acerbi and Stubbersfield(2023)}]{acerbi2023large}
Acerbi, A.; and Stubbersfield, J.~M. 2023.
\newblock Large language models show human-like content biases in transmission
  chain experiments.
\newblock \emph{Proceedings of the National Academy of Sciences}, 120(44):
  e2313790120.

\bibitem[{Achiam et~al.(2023)Achiam, Adler, Agarwal, Ahmad, Akkaya, Aleman,
  Almeida, Altenschmidt, Altman, Anadkat et~al.}]{achiam2023gpt}
Achiam, J.; Adler, S.; Agarwal, S.; Ahmad, L.; Akkaya, I.; Aleman, F.~L.;
  Almeida, D.; Altenschmidt, J.; Altman, S.; Anadkat, S.; et~al. 2023.
\newblock GPT-4 Technical Report.
\newblock \emph{arXiv preprint arXiv:2303.08774}.

\bibitem[{Ahn et~al.(2022)Ahn, Brohan, Brown, Chebotar, Cortes, David, Finn,
  Fu, Gopalakrishnan, Hausman, Herzog, Ho, Hsu, Ibarz, Ichter, Irpan, Jang,
  Ruano, Jeffrey, Jesmonth, Joshi, Julian, Kalashnikov, Kuang, Lee, Levine, Lu,
  Luu, Parada, Pastor, Quiambao, Rao, Rettinghouse, Reyes, Sermanet, Sievers,
  Tan, Toshev, Vanhoucke, Xia, Xiao, Xu, Xu, Yan, and Zeng}]{saycan-ahn2022i}
Ahn, M.; Brohan, A.; Brown, N.; Chebotar, Y.; Cortes, O.; David, B.; Finn, C.;
  Fu, C.; Gopalakrishnan, K.; Hausman, K.; Herzog, A.; Ho, D.; Hsu, J.; Ibarz,
  J.; Ichter, B.; Irpan, A.; Jang, E.; Ruano, R.~J.; Jeffrey, K.; Jesmonth, S.;
  Joshi, N.~J.; Julian, R.; Kalashnikov, D.; Kuang, Y.; Lee, K.-H.; Levine, S.;
  Lu, Y.; Luu, L.; Parada, C.; Pastor, P.; Quiambao, J.; Rao, K.; Rettinghouse,
  J.; Reyes, D.; Sermanet, P.; Sievers, N.; Tan, C.; Toshev, A.; Vanhoucke, V.;
  Xia, F.; Xiao, T.; Xu, P.; Xu, S.; Yan, M.; and Zeng, A. 2022.
\newblock Do As I Can, Not As I Say: Grounding Language in Robotic Affordances.
\newblock arXiv:2204.01691.

\bibitem[{Akalin and Loutfi(2021)}]{akalin2021reinforcement}
Akalin, N.; and Loutfi, A. 2021.
\newblock Reinforcement learning approaches in social robotics.
\newblock \emph{Sensors}, 21(4): 1292.

\bibitem[{Bertacchini et~al.(2023)Bertacchini, Demarco, Scuro, Pantano, and
  Bilotta}]{bertacchini2023social}
Bertacchini, F.; Demarco, F.; Scuro, C.; Pantano, P.; and Bilotta, E. 2023.
\newblock A social robot connected with chatGPT to improve cognitive
  functioning in ASD subjects.
\newblock \emph{Frontiers in Psychology}, 14.

\bibitem[{Billing, Ros{\'e}n, and Lamb(2023)}]{billing2023language}
Billing, E.; Ros{\'e}n, J.; and Lamb, M. 2023.
\newblock Language models for human-robot interaction.
\newblock In \emph{ACM/IEEE International Conference on Human-Robot
  Interaction, March 13--16, 2023, Stockholm, Sweden}, 905--906. ACM Digital
  Library.

\bibitem[{Cabibihan et~al.(2013)Cabibihan, Javed, Ang, and
  Aljunied}]{cabibihan2013robots}
Cabibihan, J.-J.; Javed, H.; Ang, M.; and Aljunied, S.~M. 2013.
\newblock Why robots? A survey on the roles and benefits of social robots in
  the therapy of children with autism.
\newblock \emph{International journal of social robotics}, 5: 593--618.

\bibitem[{Clabaugh and Matari{\'c}(2019)}]{clabaugh2019escaping}
Clabaugh, C.; and Matari{\'c}, M. 2019.
\newblock Escaping oz: Autonomy in socially assistive robotics.
\newblock \emph{Annual Review of Control, Robotics, and Autonomous Systems}, 2:
  33--61.

\bibitem[{Cui et~al.(2024)Cui, Ma, Cao, Ye, Zhou, Liang, Chen, Lu, Yang, Liao
  et~al.}]{cui2024survey}
Cui, C.; Ma, Y.; Cao, X.; Ye, W.; Zhou, Y.; Liang, K.; Chen, J.; Lu, J.; Yang,
  Z.; Liao, K.-D.; et~al. 2024.
\newblock A survey on multimodal large language models for autonomous driving.
\newblock In \emph{Proceedings of the IEEE/CVF Winter Conference on
  Applications of Computer Vision}, 958--979.

\bibitem[{Erich, Hirokawa, and Suzuki(2017)}]{erich2017systematic}
Erich, F.; Hirokawa, M.; and Suzuki, K. 2017.
\newblock A systematic literature review of experiments in socially assistive
  robotics using humanoid robots.
\newblock \emph{arXiv preprint arXiv:1711.05379}.

\bibitem[{Feil-Seifer and Matari{\'c}(2011)}]{feil2011socially}
Feil-Seifer, D.; and Matari{\'c}, M.~J. 2011.
\newblock Socially assistive robotics.
\newblock \emph{IEEE Robotics \& Automation Magazine}, 18(1): 24--31.

\bibitem[{Gandhi et~al.(2023)Gandhi, Fr{\"a}nken, Gerstenberg, and
  Goodman}]{gandhi2023understanding}
Gandhi, K.; Fr{\"a}nken, J.-P.; Gerstenberg, T.; and Goodman, N.~D. 2023.
\newblock Understanding social reasoning in language models with language
  models.
\newblock \emph{arXiv preprint arXiv:2306.15448}.

\bibitem[{Ge et~al.(2023)Ge, Chen, Chen, Chen, Chen, Yan, Zhu, Lin, Xie, Wang
  et~al.}]{ge2023mllm}
Ge, W.; Chen, S.; Chen, G.; Chen, J.; Chen, Z.; Yan, S.; Zhu, C.; Lin, Z.; Xie,
  W.; Wang, X.; et~al. 2023.
\newblock MLLM-Bench, Evaluating Multi-modal LLMs using GPT-4V.
\newblock \emph{arXiv preprint arXiv:2311.13951}.

\bibitem[{Grassi, Recchiuto, and Sgorbissa(2022)}]{grassi2022knowledge}
Grassi, L.; Recchiuto, C.~T.; and Sgorbissa, A. 2022.
\newblock Knowledge-grounded dialogue flow management for social robots and
  conversational agents.
\newblock \emph{International Journal of Social Robotics}, 14(5): 1273--1293.

\bibitem[{Huang et~al.(2023)Huang, Ruan, Huang, Jin, Dong, Wu, Bensalem, Mu,
  Qi, Zhao et~al.}]{huang2023survey}
Huang, X.; Ruan, W.; Huang, W.; Jin, G.; Dong, Y.; Wu, C.; Bensalem, S.; Mu,
  R.; Qi, Y.; Zhao, X.; et~al. 2023.
\newblock A Survey of Safety and Trustworthiness of Large Language Models
  through the Lens of Verification and Validation.
\newblock \emph{arXiv preprint arXiv:2305.11391}.

\bibitem[{Irfan, Kuoppam{\"a}ki, and Skantze(2023)}]{irfan2023between}
Irfan, B.; Kuoppam{\"a}ki, S.-M.; and Skantze, G. 2023.
\newblock Between Reality and Delusion: Challenges of Applying Large Language
  Models to Companion Robots for Open-Domain Dialogues with Older Adults.

\bibitem[{Jia et~al.(2021)Jia, Yang, Xia, Chen, Parekh, Pham, Le, Sung, Li, and
  Duerig}]{jia2021scaling}
Jia, C.; Yang, Y.; Xia, Y.; Chen, Y.-T.; Parekh, Z.; Pham, H.; Le, Q.; Sung,
  Y.-H.; Li, Z.; and Duerig, T. 2021.
\newblock Scaling up visual and vision-language representation learning with
  noisy text supervision.
\newblock In \emph{International conference on machine learning}, 4904--4916.
  PMLR.

\bibitem[{Koh, Ang, and Casey(2021)}]{koh2021impacts}
Koh, W.~Q.; Ang, F. X.~H.; and Casey, D. 2021.
\newblock Impacts of low-cost robotic pets for older adults and people with
  dementia: scoping review.
\newblock \emph{JMIR rehabilitation and assistive technologies}, 8(1): e25340.

\bibitem[{Koh et~al.(2022)Koh, Whelan, Heins, Casey, Toomey, and
  Dr{\"o}es}]{koh2022usability}
Koh, W.~Q.; Whelan, S.; Heins, P.; Casey, D.; Toomey, E.; and Dr{\"o}es, R.-M.
  2022.
\newblock The usability and impact of a low-cost pet robot for older adults and
  people with dementia: qualitative content analysis of user experiences and
  perceptions on consumer websites.
\newblock \emph{JMIR aging}, 5(1): e29224.

\bibitem[{Lee et~al.(2023)Lee, Jung, Kang, and Hahn}]{lee2023developing}
Lee, Y.~K.; Jung, Y.; Kang, G.; and Hahn, S. 2023.
\newblock Developing Social Robots with Empathetic Non-Verbal Cues Using Large
  Language Models.
\newblock \emph{arXiv preprint arXiv:2308.16529}.

\bibitem[{Li et~al.(2021)Li, Mu, Sun, Song, Su, and Zhang}]{li2021intention}
Li, Z.; Mu, Y.; Sun, Z.; Song, S.; Su, J.; and Zhang, J. 2021.
\newblock Intention understanding in human--robot interaction based on
  visual-NLP semantics.
\newblock \emph{Frontiers in Neurorobotics}, 14: 610139.

\bibitem[{Lim et~al.(2023)Lim, Sa, MacDonald, and Ahn}]{lim2023sign}
Lim, J.; Sa, I.; MacDonald, B.; and Ahn, H.~S. 2023.
\newblock A Sign Language Recognition System with Pepper,
  Lightweight-Transformer, and LLM.
\newblock \emph{arXiv preprint arXiv:2309.16898}.

\bibitem[{Liyanage et~al.(2020)Liyanage, Kuziemsky, Terry, Schreiber,
  Jonnagaddala, de~Lusignan, McGovern, Hinton, Corra, Munro
  et~al.}]{liyanage2020ethical}
Liyanage, H.; Kuziemsky, C.; Terry, A.; Schreiber, R.; Jonnagaddala, J.;
  de~Lusignan, S.; McGovern, A.; Hinton, W.; Corra, A.; Munro, N.; et~al. 2020.
\newblock Ethical use of electronic health record data and artificial
  intelligence: recommendations of the primary care informatics working group
  of the international medical informatics association.
\newblock \emph{Yearbook of medical informatics}, 29(01): 051--057.

\bibitem[{Matari{\'c} and Scassellati(2016)}]{mataric2016socially}
Matari{\'c}, M.~J.; and Scassellati, B. 2016.
\newblock Socially assistive robotics.
\newblock \emph{Springer handbook of robotics}, 1973--1994.

\bibitem[{Mishra et~al.(2023)Mishra, Verdonschot, Hagoort, and
  Skantze}]{mishra2023real}
Mishra, C.; Verdonschot, R.; Hagoort, P.; and Skantze, G. 2023.
\newblock Real-time emotion generation in human-robot dialogue using large
  language models.
\newblock \emph{Frontiers in Robotics and AI}, 10.

\bibitem[{Pinto-Bernal et~al.(2022)Pinto-Bernal, Cespedes, Castro, Munera, and
  Cifuentes}]{pinto2022physical}
Pinto-Bernal, M.~J.; Cespedes, N.; Castro, P.; Munera, M.; and Cifuentes, C.~A.
  2022.
\newblock Physical human-robot interaction influence in ASD therapy through an
  affordable soft social robot.
\newblock \emph{Journal of Intelligent \& Robotic Systems}, 105(3): 67.

\bibitem[{Radford et~al.(2021)Radford, Kim, Hallacy, Ramesh, Goh, Agarwal,
  Sastry, Askell, Mishkin, Clark et~al.}]{radford2021learning}
Radford, A.; Kim, J.~W.; Hallacy, C.; Ramesh, A.; Goh, G.; Agarwal, S.; Sastry,
  G.; Askell, A.; Mishkin, P.; Clark, J.; et~al. 2021.
\newblock Learning transferable visual models from natural language
  supervision.
\newblock In \emph{International conference on machine learning}, 8748--8763.
  PMLR.

\bibitem[{Randall(2019)}]{randall2019survey}
Randall, N. 2019.
\newblock A survey of robot-assisted language learning (RALL).
\newblock \emph{ACM Transactions on Human-Robot Interaction (THRI)}, 9(1):
  1--36.

\bibitem[{Robinson et~al.(2023)Robinson, Tidd, Campbell, Kuli{\'c}, and
  Corke}]{robinson2023robotic}
Robinson, N.; Tidd, B.; Campbell, D.; Kuli{\'c}, D.; and Corke, P. 2023.
\newblock Robotic vision for human-robot interaction and collaboration: A
  survey and systematic review.
\newblock \emph{ACM Transactions on Human-Robot Interaction}, 12(1): 1--66.

\bibitem[{Sap et~al.(2019)Sap, Rashkin, Chen, Le~Bras, and
  Choi}]{sap-etal-2019-social}
Sap, M.; Rashkin, H.; Chen, D.; Le~Bras, R.; and Choi, Y. 2019.
\newblock Social {IQ}a: Commonsense Reasoning about Social Interactions.
\newblock In Inui, K.; Jiang, J.; Ng, V.; and Wan, X., eds., \emph{Proceedings
  of the 2019 Conference on Empirical Methods in Natural Language Processing
  and the 9th International Joint Conference on Natural Language Processing
  (EMNLP-IJCNLP)}, 4463--4473. Hong Kong, China: Association for Computational
  Linguistics.

\bibitem[{Scoglio et~al.(2019)Scoglio, Reilly, Gorman, and
  Drebing}]{scoglio2019use}
Scoglio, A.~A.; Reilly, E.~D.; Gorman, J.~A.; and Drebing, C.~E. 2019.
\newblock Use of social robots in mental health and well-being research:
  systematic review.
\newblock \emph{Journal of medical Internet research}, 21(7): e13322.

\bibitem[{Shi et~al.(2022)Shi, Groechel, Jain, Chima, Rudovic, and
  Matari{\'c}}]{shi2022toward}
Shi, Z.; Groechel, T.~R.; Jain, S.; Chima, K.; Rudovic, O.; and Matari{\'c},
  M.~J. 2022.
\newblock Toward personalized affect-aware socially assistive robot tutors for
  long-term interventions with children with autism.
\newblock \emph{ACM Transactions on Human-Robot Interaction (THRI)}, 11(4):
  1--28.

\bibitem[{Shishehgar, Kerr, and Blake(2018)}]{shishehgar2018systematic}
Shishehgar, M.; Kerr, D.; and Blake, J. 2018.
\newblock A systematic review of research into how robotic technology can help
  older people.
\newblock \emph{Smart Health}, 7: 1--18.

\bibitem[{Singh et~al.(2023)Singh, Blukis, Mousavian, Goyal, Xu, Tremblay, Fox,
  Thomason, and Garg}]{singh2023progprompt}
Singh, I.; Blukis, V.; Mousavian, A.; Goyal, A.; Xu, D.; Tremblay, J.; Fox, D.;
  Thomason, J.; and Garg, A. 2023.
\newblock Progprompt: Generating situated robot task plans using large language
  models.
\newblock In \emph{2023 IEEE International Conference on Robotics and
  Automation (ICRA)}, 11523--11530. IEEE.

\bibitem[{Skantze(2021)}]{skantze2021turn}
Skantze, G. 2021.
\newblock Turn-taking in conversational systems and human-robot interaction: a
  review.
\newblock \emph{Computer Speech \& Language}, 67: 101178.

\bibitem[{Spitale, Axelsson, and Gunes(2023)}]{spitale2023vita}
Spitale, M.; Axelsson, M.; and Gunes, H. 2023.
\newblock VITA: A Multi-modal LLM-based System for Longitudinal, Autonomous,
  and Adaptive Robotic Mental Well-being Coaching.
\newblock \emph{arXiv preprint arXiv:2312.09740}.

\bibitem[{Wake et~al.(2023)Wake, Kanehira, Sasabuchi, Takamatsu, and
  Ikeuchi}]{wake2023gpt}
Wake, N.; Kanehira, A.; Sasabuchi, K.; Takamatsu, J.; and Ikeuchi, K. 2023.
\newblock GPT Models Meet Robotic Applications: Co-Speech Gesturing Chat
  System.
\newblock \emph{arXiv preprint arXiv:2306.01741}.

\bibitem[{Wang et~al.(2022)Wang, Lan, Liu, Ouyang, Qin, Lu, Chen, Zeng, and
  Yu}]{wang2022generalizing}
Wang, J.; Lan, C.; Liu, C.; Ouyang, Y.; Qin, T.; Lu, W.; Chen, Y.; Zeng, W.;
  and Yu, P. 2022.
\newblock Generalizing to unseen domains: A survey on domain generalization.
\newblock \emph{IEEE Transactions on Knowledge and Data Engineering}.

\bibitem[{Wu et~al.(2023)Wu, Yao, Zhang, Song, Ouyang, and
  Wang}]{wu2023gpt4vis}
Wu, W.; Yao, H.; Zhang, M.; Song, Y.; Ouyang, W.; and Wang, J. 2023.
\newblock GPT4Vis: What Can GPT-4 Do for Zero-shot Visual Recognition?
\newblock \emph{arXiv preprint arXiv:2311.15732}.

\bibitem[{Xu et~al.(2022)Xu, Alon, Neubig, and Hellendoorn}]{xu2022systematic}
Xu, F.~F.; Alon, U.; Neubig, G.; and Hellendoorn, V.~J. 2022.
\newblock A systematic evaluation of large language models of code.
\newblock In \emph{Proceedings of the 6th ACM SIGPLAN International Symposium
  on Machine Programming}, 1--10.

\bibitem[{Youssef et~al.(2022)Youssef, Said, Alkork, and
  Beyrouthy}]{youssef2022survey}
Youssef, K.; Said, S.; Alkork, S.; and Beyrouthy, T. 2022.
\newblock A survey on recent advances in social robotics.
\newblock \emph{Robotics}, 11(4): 75.

\bibitem[{Zhang et~al.(2023{\natexlab{a}})Zhang, Huang, Jin, and
  Lu}]{zhang2023vision}
Zhang, J.; Huang, J.; Jin, S.; and Lu, S. 2023{\natexlab{a}}.
\newblock Vision-language models for vision tasks: A survey.
\newblock \emph{arXiv preprint arXiv:2304.00685}.

\bibitem[{Zhang et~al.(2023{\natexlab{b}})Zhang, Li, Cui, Cai, Liu, Fu, Huang,
  Zhao, Zhang, Chen et~al.}]{zhang2023siren}
Zhang, Y.; Li, Y.; Cui, L.; Cai, D.; Liu, L.; Fu, T.; Huang, X.; Zhao, E.;
  Zhang, Y.; Chen, Y.; et~al. 2023{\natexlab{b}}.
\newblock Siren's Song in the AI Ocean: A Survey on Hallucination in Large
  Language Models.
\newblock \emph{arXiv preprint arXiv:2309.01219}.

\end{thebibliography}

\documentclass[letterpaper]{article} 
\usepackage{aaai24}  
\usepackage{times}  
\usepackage{helvet}  
\usepackage{courier}  
\usepackage[hyphens]{url}  
\usepackage{graphicx} 
\usepackage{natbib}  
\usepackage{caption} 
\usepackage{algorithm}
\usepackage{algorithmic}

\usepackage{newfloat}
\usepackage{listings}
\DeclareCaptionStyle{ruled}{labelfont=normalfont,labelsep=colon,strut=off} 
\floatstyle{ruled}
\newfloat{listing}{tb}{lst}{}
\floatname{listing}{Listing}
\title{How Can Large Language Models Enable Better Socially Assistive\\ Human-Robot Interaction: A Brief Survey}
\author{
    Zhonghao Shi, Ellen Landrum,
    Amy O'Connell,
    Mina Kian,
    Leticia Pinto-Alva,
    Kaleen Shrestha,
    Xiaoyuan Zhu,
    Maja J Matari\'c
}
\affiliations{
    University of Southern California, Los Angeles, United States\\

     \{zhonghas, elandrum, amy.dell, kian, pintoalv, kshresth, xzhu9839, mataric\}@usc.edu
}

\usepackage{bibentry}

\begin{document}

\maketitle

\begin{abstract}
Socially assistive robots (SARs) have shown great success in providing personalized cognitive-affective support for user populations with special needs such as older adults, children with autism spectrum disorder (ASD), and individuals with mental health challenges. The large body of work on SAR demonstrates its potential to provide at-home support that complements clinic-based interventions delivered by mental health professionals, making these interventions more effective and accessible. However, there are still several major technical challenges that hinder SAR-mediated interactions and interventions from reaching human-level social intelligence and efficacy. With the recent advances in large language models (LLMs), there is an increased potential for novel applications within the field of SAR that can significantly expand the current capabilities of SARs. However, incorporating LLMs  introduces new risks and ethical concerns that have not yet been encountered, and must be carefully be addressed to safely deploy these more advanced systems.
In this work, we aim to conduct a brief survey on the use of LLMs in SAR technologies, and discuss the potentials and risks of applying LLMs to the following three major technical challenges of SAR: 1) natural language dialog; 2) multimodal understanding; 3) LLMs as robot policies. 
\end{abstract}

\section{Introduction}

At the intersection of assistive robots and socially interactive robots, {\it socially assistive robots} aim to provide assistance to support users' cognitive-affective needs through social interaction~\cite{feil2011socially}. Past studies have demonstrated the benefits and potential of deploying SARs to support a diverse set of user populations, including individuals with mental health challenges~\cite{scoglio2019use}, children with ASD~\cite{cabibihan2013robots}, older adults~\cite{shishehgar2018systematic}, and K-12 students~\cite{randall2019survey}. With more affordable robot hardware~\cite{koh2021impacts, pinto2022physical, koh2022usability}, SARs have the potential to lower the socio-economic barriers that limit access to personalized therapies, companionship, and education. 
However, prior work has also demonstrated that the existing SAR interactions have not yet been able to understand multimodal (language and visual) social dynamics~\cite{robinson2023robotic, li2021intention} and respond with human-like dialog~\cite{skantze2021turn} and actions~\cite{akalin2021reinforcement}.

With the recent advances in natural language processing (NLP) research, large language models (LLMs) such as GPT-4 have shown tremendous success in tasks both within the field of NLP (such as language modeling, question answering, and translation)~\cite{achiam2023gpt} and outside (such as programming~\cite{xu2022systematic}, robot planning~\cite{singh2023progprompt}, and autonomous driving~\cite{cui2024survey}). These capabilities of LLMs may open new possibilities for tackling the core technical challenges of SAR, and help us get closer to achieving more effective, human-level social assistance for users with differences.

In this work, we categorize the core technical challenges of SAR into three areas: 1) natural language dialog; 2) multimodal user understanding; 3) LLMs as robot policies. To survey the existing work on LLM-powered SARs in these three categories of technical challenges, we used Google Scholar to search the relevant papers in major human-robot interaction conferences, journals, and arXiv. In each of the following sections, we aim to: identify the potential of LLMs for SAR, survey existing work, and discuss future directions. This research was supported by the National Science Foundation Grant ITE-2236320 and IIS-1925083.

\section{Natural Language Dialogue }

Natural language dialogue is at the core of human-centered social interaction. Yet, prior to the recent breakthroughs in LLM technologies, SARs mainly relied on Wizard-of-Oz teleoperation~\cite{erich2017systematic} or predefined rule-based dialogue management systems~\cite{erich2017systematic, youssef2022survey}. SARs employing traditional non-LLM-based conversational systems are limited by their inability to accurately interpret human dialogue, limited vocabulary in dialogue generation, lack of understanding of context and ability to personalize, and lack of ability to effectively utilize online resources~\cite{grassi2022knowledge}.




By applying state-of-art LLM models such as GPT-4~\cite{achiam2023gpt}, recent work on LLM-powered SAR has been mainly focused on enabling more accurate dialog understanding and more human-like and context-aware dialogue generation. LLM-powered SARs are able to produce varied dialogue while staying on topic ~\cite{billing2023language}. They can engage in more natural, flexible conversation with users from populations of interest, such as older adults and children with autism spectrum disorder~\cite{bertacchini2023social}. \citet{spitale2023vita} designed an LLM-powered SAR as a motivational coach with both informative and emotional objectives, demonstrating that LLMs can be used to understand long-horizon context and enable long-term personalization. In instructional settings, LLM-powered SARs combine the vast knowledge base and interactive content-delivery of LLMs with the the capabilities and engaging nature of physically embodied agents~\cite{wake2023gpt}. 

Despite this progress, \citet{irfan2023between} showed that hallucinations, obsolete information, latency, and disengagement cues may still cause user frustration and confusion, which could be detrimental to socially assistive human-robot interaction. 
More exploration is needed to overcome these limitations and harness the power of LLMs in ways that better align with the goals of SAR interactions.

\section{Multimodal User Understanding}

To enable successful socially assistive human-robot interactions, a SAR needs to understand the user's cognitive-affective state (user engagement, affect, and intent) from multimodal perceptual data (language, visual, and audio)~\cite{youssef2022survey}. Existing work on multimodal social understanding has relied on training and fine-tuning machine learning (ML) models with data collected from previous SAR deployments~\cite{robinson2023robotic}. However, the definition of cognitive-affective states may vary in different social contexts. Due to the independent and identically distributed (IID) assumption made by ML model training~\cite{wang2022generalizing}, existing ML models struggle to generalize effectively and quickly to test data that are distributed differently from the training data, particularly in the context of SAR~\cite{shi2022toward}.

Multimodal language models (MLMs) such as state-of-the-art vision-language models like CLIP~\cite{radford2021learning}, ALIGN~\cite{jia2021scaling}, and GPT-4V~\cite{achiam2023gpt}, have shown promising zero-shot performance on a variety of human-centered visual tasks~\cite{zhang2023vision, wu2023gpt4vis}. Furthermore, these MLMs also demonstrate impressive few-shot capabilities of quickly adapting via prompting with natural language~\cite{ge2023mllm}. This indicates that MLMs may also be capable of adapting to novel social context for more generalizable and accurate multimodal social understanding. Despite the recent progress in computer vision and robotics, using MLMs for SAR is still largely unexplored, but this direction of research shows great potential for significantly enabling better multimodal social understanding for socially assistive human-robot interaction.

\section{LLMs as Robot Policies}

In an ideal socially assistive human-robot interaction, a SAR should fluently learn and reason about the user's states and provide the best feedback or action as assistance. The space of user's states and robot feedback or actions can be large and continuous~\cite{clabaugh2019escaping} and existing approaches to encoding SAR policies, such as rule-based system and reinforcement learning, are not efficiently trained or sufficiently robust on large and continuous spaces with limited amounts of data~\cite{akalin2021reinforcement}. Past work has often circumvented this problem by constraining interactions to pre-defined tablet/computer games with small user state and action spaces~\cite{clabaugh2019escaping}. LLMs may help to relax this constraint and enable more natural interaction by allowing continuous, more human-like formulation of space for user states and robot actions.

Recent studies in robotics and NLP have successfully employed LLMs as robot policies in the setting including autonomous driving~\cite{cui2024survey}, robot task planning~\cite{singh2023progprompt, saycan-ahn2022i}, social common sense~\cite{sap-etal-2019-social} and social reasoning~\cite{gandhi2023understanding}. In the field of SAR, existing work has mainly focused on applying LLMs to matching the affective state of robot feedback with the sentiment of the robot's dialog~\cite{lee2023developing, mishra2023real, lim2023sign}. These studies have shown that LLM-powered robot policies for gesture matching enable more context-aware and natural robot gesture. Research using LLMs as robot policies has not yet explored 1) how to enable SARs to form robot policies for spontaneous tasks instead of only pre-defined ones, such as helping children with ASD navigate through an unpleasant social interaction they just encountered during their school day; 2) how to enable SAR policies to engage users with educational tasks while keeping them both challenged and encouraged; 3) how to reason socially about the user's intent and needs with partially observable information based on multimodal data; and 4) how to enable personalized SAR policies to quickly align with each user's unique needs, personality, and values, so the SAR can be more helpful and empathetic to each user.





\section{Risks and Safety Considerations}

Because SAR aims to support vulnerable populations including children and individuals with physical and/or mental health challenges~\cite{mataric2016socially}, it is crucial to ensure absolute safety during socially assistive human-robot interactions. Despite the great success in LLMs, their lack of explainability and theoretical safety guarantees~\cite{huang2023survey} may introduce significant risks and concerns by 1) amplifying unfairness and human bias~\cite{acerbi2023large}; 2) harming data security and privacy by unethical use of personal data~\cite{liyanage2020ethical}; and 3) hallucination behaviors causing potential harms to users~\cite{zhang2023siren}. For these reasons, the trustworthiness of LLM-powered SAR needs to be extensively evaluated before autonomous system can be deployed with vulnerable populations without human monitoring and oversight. As the first survey paper on this topic, this work aims to inform and stimulate research toward leveraging the tremendous potential of LLM-powered SARs and address the risks and safety concerns.

\bibliography{aaai24}

\end{document}